OpenStudy (anonymous):

Jack wants to find the solution to a system of equations using y = x + 3 and y = x^2 + 6x - 21. Betty says that Jack can solve x^2 + 5x - 24 = 0 to find the x-coordinates of the solutions of the system. Explain and demonstrate why Betty is correct.

OpenStudy (whpalmer4):

A solution to that system of equations will be a point \((x,y)\) where both equations are true. If they are true at the same point, then \(x\) in one equation equals \(x\) in the other equation, and similarly \(y\) in one equation equals \(y\) in the other. That means we can say \[y = y\]\[x+3 = x^2+6x-21\] What do you get when you collect all of the like terms in that equation?

OpenStudy (anonymous):

-5x= x^2 - 24?